# An Empirical Analysis of Compute-Optimal Inference for Problem-Solving with Language Models

### v1

@@ Abstract
The optimal training configurations of large language models (LLMs) with respect to model sizes and compute budgets have been extensively studied. But how to optimally configure LLMs during inference has not been explored in sufficient depth. We study compute-optimal inference: designing models and inference strategies that optimally trade off additional inference-time compute for improved performance. As a first step towards understanding and designing compute-optimal inference methods, we assessed the effectiveness and computational efficiency of multiple inference strategies such as Greedy Search, Majority Voting, Best-of-N, Weighted Voting, and their variants on two different Tree Search algorithms, involving different model sizes and computational budgets. We found that a smaller language model with a novel tree search algorithm typically achieves a Pareto-optimal trade-off. These results highlight the potential benefits of deploying smaller models equipped with more sophisticated decoding algorithms in budget-constrained scenarios, e.g., on end-devices, to enhance problem-solving accuracy. For instance, we show that the Llemma-7B model can achieve competitive accuracy to a Llemma-34B model on MATH500 while using 2×2$\times$2 × less FLOPs. Our findings could potentially apply to any generation task with a well-defined measure of success.
@@ 1 Introduction
Scaling laws of neural networks (Hestness et al., 2017; Rosenfeld et al., 2019) have been established across a range of domains, including language modeling (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2023), image modeling (Henighan et al., 2020; Yu et al., 2022; Peebles and Xie, 2023), video modeling (Brooks et al., 2024), reward modeling (Gao et al., 2023), and board games (Jones, 2021). These studies have demonstrated how model performance is influenced by both the size of the model and the amount of training computation. However, there is limited knowledge on how varying the compute during inference affects model performance after the model has been trained.
Figure 1: The inference computation scaling laws of Pythia exhibited in error rate on the GSM8K test set. We evaluate Pythia model using various sizes and various numbers of sampled solutions for majority voting. The left panel shows the error rate for each model size decreases steadily when the computation increases and converges at the end. The right panel shows the model performances given inference FLOPs budgets. In particular, the three stars highlight the optimal model size under $\color[rgb]{0.75390625,0.92578125,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.75390625,0.92578125,0.578125}2^{41}$, $\color[rgb]{0.3515625,0.8828125,0.85546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.3515625,0.8828125,0.85546875}2^{44}$, and $\color[rgb]{0.4140625,0.171875,0.984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4140625,0.171875,0.984375}2^{47}$ FLOPs, indicating that the optimal model size can vary given different budgets. Both the $x$ and $y$ axes are shown in $\log$ scale.
@@
To improve the task performance of large language models (LLMs), inference techniques typically involve additional computation as a performance maximization step at inference time (Nye et al., 2021; Wei et al., 2022; Wang et al., 2022b; Yao et al., 2023; Chen et al., 2024b). This cost must be taken into account for compute-optimal inference. For example, a Monte Carlo Tree Search (MCTS) method  (Jones, 2021) may improve task performance, but potentially require much more compute than simply sampling solutions multiple times. Generally speaking, we need a comprehensive understanding of how various inference-time methods (e.g., Best-of-N, Majority Voting) trade off between performance and cost. To improve our understanding, this paper presents a thorough empirical evaluation with careful analysis over various configurations of representative LLMs and inference algorithms.
@@
Specifically, we explore how to select an optimal size for the language model and an effective inference strategy (e.g., Greedy Search, Majority Voting, Best-of-N, Weighted Voting, and their Tree Search variants) to maximize performance (i.e., accuracy) with a given compute budget. We control the inference computation (FLOPs) of a fixed model by utilizing more tokens through the language model [^1] Following Uesato et al. (2022), we refer to the main language model generating outputs as the policy model. It can be paired with a reward model, which scores outputs from the policy model to facilitate inference., sampling further candidate solutions, and ranking them with a reward model. In this study, we analyze the performances of fine-tuned models of various sizes given different inference FLOPs on mathematical reasoning benchmarks (e.g., GSM8K test set (Cobbe et al., 2021a) and MATH500 test set (Hendrycks et al., 2021b; Lightman et al., 2023b)). Our experiments cover several model families, including general-purpose LLMs (e.g., Pythia (Biderman et al., 2023) & Mistral (Jiang et al., 2023)) as well as math-specialized ones (e.g., Llemma (Azerbayev et al., 2023)).
@@
Our results on Pythia (Fig. 1) illustrate the scaling effect of inference compute budget across various model sizes. Typically, increasing computation leads to higher accuracy until the accuracy reaches saturation. As the computation budget increases, the smaller model initially performs better than the larger ones. Once the accuracy of the smaller models saturates, the larger models begin to show their capability. The right panel of Figure 1 demonstrates that the optimal model size for inference varies with different levels of computation. However, in real-world deployment, the available computation is typically much lower than the point where the accuracy of relatively small models saturates and larger models begin to show their advantage (as shown in Figure 4, where the 7B model outperforms the 34B model until 128 Llemma 7B solutions are sampled). This indicates that relatively smaller models could be more compute-optimal for inference.
@@
We have also found that the commonly-used MCTS method does not perform well with weighted voting, as it often yields many unfinished solutions, hence having less effective votes. To address this issue, we propose a novel tree search algorithm, REward BAlanced SEarch (REBASE), which pairs well with weighted voting and improves the Pareto-optimal trade-off between accuracy and inference compute. The key idea of REBASE is to use a node-quality based reward to control the exploitation and pruning properties of tree search, while ensuring enough candidate solutions for voting or selection.
@@
In our experiments, REBASE consistently outperforms sampling and MCTS methods across all settings, models, and tasks. Importantly, we find that REBASE with a smaller language model typically achieves a Pareto-optimal trade-off. For instance, we show that the Llemma-7B model can achieve competitive accuracy to a Llemma-34B model while using 2×2$\times$2 × less FLOPs when evaluating on MATH500 (Fig. 4) or GSM8K (Fig. 5). These findings underscore the advantages of using smaller models with advanced inference-time algorithms on end-devices to improve problem-solving.
@@ 2 Related Works
Large language models have made significant progress in recent years, and have exhibited strong reasoning abilities (Brown et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2022; Lewkowycz et al., 2022). Mathematical problem solving is a key task for measuring LLM reasoning abilities (Cobbe et al., 2021a; Hendrycks et al., 2021b). (Ling et al., 2017) first developed the method of producing step by step solutions that lead to the final answer. Later, (Cobbe et al., 2021b) extended the work by finetuning the pre-trained language model on a large dataset to solve math word problems, a verifier is trained for evaluating solutions and ranking solutions. Nye et al. (2021) train models to use a scratchpad and improve their performance on algorithmic tasks. Wei et al. (2022) demonstrate that the reasoning ability of a language model can be elicited through the prompting. Subsequent research (Kojima et al., 2022; Lewkowycz et al., 2022; Zhou et al., 2022) in reasoning tasks has also highlighted the efficacy of rationale augmentation. We choose problem solving in mathematics as the task to study the compute-optimal strategy since it allows us to accurately evaluate the problem solving ability of LLMs.
@@
A variety of inference (also called decoding) strategies have been developed to generate sequences with a trained model. Deterministic methods such as greedy decoding and beam search (Teller, 2000; Graves, 2012) find highly probable sequences, often yielding high quality results but without diversity. Sampling algorithms (e.g., temperature sampling (Ackley et al., 1985)) can produce a diverse set of results which are then aggregated to achieve higher accuracy (e.g., via Majority Voting (Wang et al., 2022a)). Recent methods combine search algorithms with modern LLMs, including breadth-first or depth-first search (Yao et al., 2023), Monte-Carlo Tree Search (MCTS) (Zhang et al., 2023; Zhou et al., 2023; Liu et al., 2024; Choi et al., 2023), and Self-evaluation Guided Beam Search (Xie et al., 2023). All of these methods show that using search at inference time can lead to performance gains in various tasks.. However, the trade-off for the improved performance is the use of compute to perform the search. Analyzing the trade-off between compute budget and LLM inference performance remains understudied. In this paper, we systematically analyze the trade-off between compute budget and problem-solving performance, and propose a tree search method that is empirically Pareto-optimal. .
@@
Process reward models (PRMs) have emerged as a technique to improve the reasoning and problem-solving capabilities of LLMs. These models assign rewards to the intermediate steps of the LLM generated sequences. PRMs have been shown effective in selecting reasoning traces with a low error rate, and for providing rewards in reinforcement learning-style algorithms (Uesato et al., 2022; Polu and Sutskever, 2020; Gudibande et al., 2023). Ma et al. (2023) applies the PRM to give rewards on the intermediate steps and guide the multi-step reasoning process. The PRM can be either trained on human labeled data (Lightman et al., 2023a) or model-labeled synthetic data (Wang et al., 2023). In our work, we use the PRM as the reward model for selecting generated solutions, and for selecting which partial solutions to explore in tree search.
Figure 2: Illustration of compute-optimal scaling laws in training and inference. The Chinchilla scaling law (Hoffmann et al., 2022) shows how to choose a model size and number of training tokens under a training-compute budget, while our work shows how to choose a model size and an inference strategy under a inference-compute budget.
@@ 3 Compute-Optimal Inference for Problem-Solving
3.1 Problem Formulation
We explore the following question: Given a fixed FLOPs budget, how should one select an optimal model size for the policy model, and an effective inference strategy to maximize performance (i.e., accuracy)? We are the first to formulate this problem and study the inference time scaling law, setting our work apart from previous scaling law studies (Fig. 2).
@@ 3.1 Problem Formulation
Here, the inference computation (FLOPs) for a fixed model can be modulated by generating more tokens with the policy model, e.g., by sampling additional candidate solutions and subsequently ranking them using a reward model. We primarily consider sampling and tree-search approaches with reranking or Majority Voting as the means to consume more tokens, including Greedy Search, Majority Voting, Best-of-N, Weighted Voting, and their variants on tree search methods.
@@ 3.2 Inference Strategies
3.2.1 Sampling-based Methods
We consider the following sampling-based inference strategies which are popularly used in practice:
  • Greedy Search. This strategy generates tokens one at a time by selecting the highest probability token at each step, without considering future steps. It is computationally efficient but often suboptimal in terms of diversity.
  • Best-of-N. This strategy, also known as rejection sampling, samples multiple solutions and chooses the one with the highest score given by the reward model.
  • Majority Voting. In this strategy, multiple model outputs are generated, and the final answer to the problem is determined by the most frequently occurring answer in all the outputs.
  • Weighted Majority Voting. This strategy is a variant of Majority Voting in which the votes are weighted based on the scores given by the reward model.
Theoretical Analysis.
Before diving into more sophisticated inference algorithms (e.g., tree search), we present theoretical results on the asymptotic behavior of voting-based strategies given infinite compute in Theorems 1 & 2. Informally, we show that the accuracy of standard/Weighted Majority Voting converges with infinite samples, and the limit only depends on the distribution modeled by the language model (and the reward model). This theoretical finding is also aligned with our empirical findings shown in Sec. 4.2. The proofs are presented in Appendix A.
@@ 3.2.1 Sampling-based Methods
Theorems 1 & 2 state the convergence of the accuracy with increasing number of samples, indicating that the performance gains of using more samples will saturate for any fixed models. The limit is determined by the likelihood of generating the correct answers through all possible reasoning paths (and the likelihood should be viewed as a weighted sum for Weighted Majority Voting). This motivates us to consider inference algorithms that search for “good” reasoning paths, such as the tree-search-based variants detailed in Sec. 3.2.2 & 3.2.3.
@@ 3.2.2 Monte Carlo Tree Search (MCTS)
Monte Carlo Tree Search (MCTS) has proven effective in domains such as board games where strategic decision-making is required (Silver et al., 2016, 2017; Jones, 2021). Recent work has shown that adapting MCTS to the context of LLMs can enhance the text generation process (Zhang et al., 2023; Zhou et al., 2023; Liu et al., 2024; Choi et al., 2023; Chen et al., 2024a; Tian et al., 2024; Chen et al., 2024a). In this context, MCTS is often paired with a value model to score and guide the exploration steps. For additional background, we provide a review of MCTS in Appendix B.
Recent work in MCTS or its variants (e.g., Tree of Thoughts (Yao et al., 2023)) mainly focus on improving the performance (e.g., accuracy) on the studied tasks. However, generic comparisons of MCTS with conventional methods like Best-of-N and Majority Voting in terms of computational budget, measured in generated tokens or processing time, are either scarce or indicating inference-time issues. For example, MCTS consumes substantially more resources, often requiring dozens of times more generated tokens than simpler methods. Specifically, a significant portion of the paths in the search tree are used to estimate and select nodes, and these paths do not necessarily become a part of the final candidate solution, although MCTS ensures that the sampled solutions comprise high-quality intermediate steps. In contrast, sampling methods generate multiple solutions in parallel and independently, and all the generated sequences are included in the candidate solutions. However, the intermediate steps in these sequences are not guaranteed to be of high quality, as there is no mechanism for pruning poor steps or exploiting promising ones.
This highlights the need for developing a new tree search method that can achieve a comparable (or better) performance as MCTS, and that is computationally less costly, just like Weighted Majority Voting and best-of-N. This need motivates the development of our new method named Reward Balanced SEarch (REBASE), as is introduced next.

### v2

# Inference Scaling Laws: An Empirical Analysis of Compute-Optimal Inference for LLM Problem-Solving

@@ Abstract
While the scaling laws of large language models (LLMs) training have been extensively studied, optimal inference configurations of LLMs remain underexplored. We study inference scaling laws and compute-optimal inference, focusing on the trade-offs between model sizes and generating additional tokens with different inference strategies. As a first step towards understanding and designing compute-optimal inference methods, we studied cost-performance trade-offs for inference strategies such as greedy search, majority voting, best-of-$n$, weighted voting, and two different tree search algorithms, using different model sizes and compute budgets. Our findings indicate smaller models (e.g., Llemma-7B) can outperform larger models given the same computation budgets, and that smaller models paired with advanced inference algorithms yield Pareto-optimal cost-performance trade-offs. For instance, the Llemma-7B model, equipped with our novel tree search algorithm, consistently outperforms Llemma-34B with standard majority voting on the MATH benchmark across all FLOPs budgets. We hope these findings contribute to a broader understanding of inference scaling laws for LLMs.
Figure 1: Inference scaling laws exhibited for Pythia (Biderman et al., 2023) models and GSM8K test error. We evaluate the error rate (lower is better) of models using various sizes and numbers of sampled solutions for weighted majority voting. Left: the error rate for each model size decreases steadily as inference-compute increases, and converges at the end. Right: the optimal model size (shown as stars for $\color[rgb]{0.84375,0.859375,0.48828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.84375,0.859375,0.48828125}2^{41}$, $\color[rgb]{0.37890625,0.98046875,0.76953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.37890625,0.98046875,0.76953125}2^{44}$, and $\color[rgb]{0.140625,0.5390625,0.9609375}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.5390625,0.9609375}2^{47}$ FLOPs) varies based on the inference-time compute budget. For instance, smaller models are compute-optimal at $\color[rgb]{0.84375,0.859375,0.48828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.84375,0.859375,0.48828125}2^{41}$ and $\color[rgb]{0.37890625,0.98046875,0.76953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.37890625,0.98046875,0.76953125}2^{44}$ FLOPs. Both axes are $\log$ scale.
@@ 1 Introduction
Scaling laws of neural networks (Hestness et al., 2017; Rosenfeld et al., 2019) have been established across a range of domains, including language modeling (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2023), image modeling (Henighan et al., 2020; Yu et al., 2022; Peebles & Xie, 2023), video modeling (Brooks et al., 2024), reward modeling (Gao et al., 2023), and board games (Jones, 2021). These studies have demonstrated how model performance is influenced by both the size of the model and the amount of training computation. However, there is limited knowledge on how varying the compute during inference affects model performance after the model has been trained.
@@
To improve the task performance of large language models (LLMs), inference techniques typically involve additional computation as a performance maximization step at inference time (Nye et al., 2021; Wei et al., 2022; Wang et al., 2022b; Yao et al., 2023; Chen et al., 2024b). The computational cost of these techniques must be taken into account for compute-optimal inference. For example, a Monte Carlo Tree Search (MCTS) method (Jones, 2021) may improve task performance, but potentially require much more compute than simply sampling solutions multiple times. Generally speaking, we need a comprehensive understanding of how various inference-time methods (e.g., best-of-$n$, majority voting (Wang et al., 2022a; Li et al., 2023)) trade off between performance and cost. To improve our understanding, this paper presents a thorough empirical evaluation with careful analysis over various configurations of representative LLMs and inference algorithms.
@@
Specifically, we explore how to select an optimal size for the language model and an effective inference strategy (e.g., greedy search, majority voting, best-of-$n$, weighted voting, and their tree-search variants) to maximize performance (i.e., accuracy) with a given compute budget. We control the inference computation (FLOPs) of a fixed model by generating more tokens through the language model [^1] Following Uesato et al. (2022), we refer to the main language model generating outputs as the policy model. It can be paired with a reward model, which scores outputs from the policy model to facilitate inference., sampling further candidate solutions, and ranking them with a reward model. We analyze the performance of fine-tuned models of various sizes given different inference FLOPs on mathematical reasoning benchmarks (e.g., GSM8K test set (Cobbe et al., 2021a) and MATH500 test set (Hendrycks et al., 2021b; Lightman et al., 2023b)). Our experiments cover several model families, including general-purpose LLMs (e.g., Pythia (Biderman et al., 2023) & Mistral (Jiang et al., 2023)) as well as math-specialized ones (e.g., Llemma (Azerbayev et al., 2023)).
@@
Our results on Pythia (Fig. 1) illustrate how performance scales with increased inference compute across various model sizes. Typically, increasing the compute budget leads to higher accuracy until the accuracy reaches saturation. As the compute budget increases, smaller models initially perform better than larger ones, but once the accuracy of the smaller models saturates, the larger models have favorable performance. The right panel of Figure 1 demonstrates that the optimal model size for inference varies with different levels of computation. However, in real-world deployment, the available computation is typically much lower than the point where the accuracy of relatively small models saturates and larger models begin to show their advantage (as shown in Figure 4, where the 7B model outperforms the 34B model until 128 Llemma 7B solutions are sampled). This indicates that relatively smaller models could be more compute-optimal for inference.
We analyse the asymptotic behaviour of sampling and voting-based inference strategies, showing their convergence upper bound and rate of convergence. Given a dataset, the accuracy of the language model will ultimately saturate to a fixed limit which is determined by the output probabilities assigned by the model, exhibiting exponential convergence speed through sampling and voting. This implies that, without an oracle verifier, simple strategies like sampling cannot achieve perfect accuracy even with an infinite number of samples, leading to diminishing returns. Therefore, this highlights the necessity for more sophisticated inference algorithms.
@@
We have also found that the commonly-used MCTS method does not perform well with weighted voting, as it often yields many unfinished solutions, hence having less effective votes. To address this issue, we propose a novel tree search algorithm, REward BAlanced SEarch (Rebase), which pairs well with weighted voting and achieves a Pareto-optimal trade-off between accuracy and inference compute. The key idea of Rebase is to use a node-quality reward to control node expansion, which eliminates the need for explicit rollouts while ensuring enough candidate solutions for voting.
@@
In our experiments, Rebase consistently outperforms sampling and MCTS methods across all settings, models, and tasks. Importantly, we find that Rebase with a smaller language model typically achieves a Pareto-optimal trade-off. For instance, we show that the Llemma-7B model can achieve competitive accuracy to a Llemma-34B model while using 2×2$\times$2 × less FLOPs when evaluating on MATH500 (Fig. 4) or GSM8K (Fig. 5). Moreover, Llemma-7B with Rebase outperforms Llemma-34B with standard majority voting across all compute budgets. Our results show the value of using smaller models with advanced inference-time algorithms, and the benefits of new algorithms for achieving better returns on inference-time compute.
Our contributions are summarized as follows:
  • We explore new inference scaling laws by evaluating the performance of various model sizes under a fixed inference strategy. We show that smaller models can outperform larger ones under the same compute budget by increasing the number of samples.
  • We provide new theoretical analysis of the scaling behavior of voting methods, presenting convergence bounds and rates. Our analysis shows performance limits and diminishing returns from sampling, pointing to the need for more sophisticated inference algorithms.
  • We formulate a new compute-optimal inference problem and propose a novel tree search algorithm, Rebase, which is compute-optimal compared to widely-used sampling and MCTS methods. Our results show benefits of using smaller models with advanced inference algorithms, and new algorithms for achieving better cost-performance tradeoffs.
@@ 2 Related Works
Large language models have made significant progress in recent years, and have exhibited strong reasoning abilities (Brown et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2022; Lewkowycz et al., 2022). Mathematical problem solving is a key task for measuring LLM reasoning abilities (Cobbe et al., 2021a; Hendrycks et al., 2021b). Ling et al. (2017) first developed the method of producing step by step solutions that lead to the final answer. Later, (Cobbe et al., 2021b) extended the work by training a verifier for evaluating and ranking solutions. Subsequent research (e.g., Lewkowycz et al. (2022)) has shown the performance benefits of inference-time techniques such as majority voting (Wang et al., 2022a) and weighted majority voting (Li et al., 2023). We choose problem solving in mathematics as the task to study compute-optimal strategies since it allows us to accurately evaluate problem solving ability.
@@
A variety of inference strategies have been developed to generate sequences with a trained model. Deterministic methods such as greedy decoding and beam search (Teller, 2000; Graves, 2012) find highly probable sequences, often yielding high quality results but without diversity. Sampling algorithms (e.g., temperature sampling (Ackley et al., 1985)) can produce a diverse set of results which are then aggregated to achieve higher accuracy (e.g., via majority voting (Wang et al., 2022a)). Recent methods combine search algorithms with LLMs, including breadth-first or depth-first search (Yao et al., 2023), Monte-Carlo Tree Search (MCTS) (Zhang et al., 2023; Zhou et al., 2023; Liu et al., 2024; Choi et al., 2023), and guided beam search (Xie et al., 2023). All of these methods show that using search at inference time can lead to performance gains in various tasks. However, the trade-off for the improved performance is the use of compute to perform the search. Analyzing the resulting cost-performance trade-offs remains understudied. In this paper, we systematically analyze the trade-off between compute budget and problem-solving performance, and propose a tree search method that is empirically Pareto-optimal.
@@
Process reward models (PRMs) have emerged as a technique to improve the reasoning and problem-solving capabilities of LLMs. These models assign rewards to the intermediate steps of the LLM generated sequences. PRMs have been shown effective in selecting reasoning traces with a low error rate, and for providing rewards in reinforcement learning-style algorithms (Uesato et al., 2022; Polu & Sutskever, 2020; Gudibande et al., 2023). Ma et al. (2023) applies a PRM to give rewards on the intermediate steps and guide the multi-step reasoning process. The PRM can be either trained on human labeled data (Lightman et al., 2023a) or model-labeled synthetic data (Wang et al., 2023). In our work, we use the PRM as the reward model for selecting generated solutions, and for selecting which partial solutions to explore in tree search.
Figure 2: Illustration of compute-optimal scaling laws in training and inference. The Chinchilla scaling law (Hoffmann et al., 2022) shows how to choose a model size and number of training tokens under a training-compute budget, while our work shows how to choose a model size and an inference strategy under an inference-compute budget.
@@ 3 Compute-Optimal Inference for Problem-Solving
We explore the following question: Given a fixed FLOPs budget, how should one select an optimal model size for the policy model, and an effective inference strategy to maximize performance (i.e., accuracy)? We are the first to formulate this problem and study the associated inference-time scaling laws, setting our work apart from previous scaling law studies (Fig. 2).
@@
Here, the inference computation (FLOPs) for a fixed model can be modulated by generating more tokens with the policy model and an inference strategy, e.g., sampling additional candidate solutions and subsequently ranking them using a reward model. As the inference strategies, we primarily consider sampling and tree-search approaches paired with re-ranking or majority voting. This includes greedy search, majority voting, best-of-$n$, weighted voting, and their tree-search variants.
@@ 3.1 Inference Strategies
We consider the following inference strategies which are popularly used in practice:
  • Greedy search. This strategy generates tokens one at a time by selecting the highest probability token at each step. It is computationally efficient but often suboptimal in terms of diversity.
  • Best-of-$n$. This strategy, also known as rejection sampling, generates a set of candidates and chooses the one with the highest score given by the reward model.
  • Majority voting. In this strategy, a set of candidates are generated, and the final answer to the problem is determined by the most frequently occurring answer in all the outputs.
  • Weighted majority voting. This strategy is a variant of majority voting in which the candidates are weighted based on the scores given by the reward model.
We say a strategy is sampling-based if it uses a standard autoregressive sampling algorithm (e.g., temperature sampling) to generate the candidate set (greedy search is separate, in that it only has a single deterministic candidate). A tree-search variant uses a tree search to generate the candidate set. Before discussing tree-search methods, we analyze sampling-based voting below.
Theoretical analysis of sampling-based voting.
We present theoretical results on the asymptotic behavior of voting-based strategies given infinite compute in Theorems 1 & 2. Informally, we show that the accuracy of standard/weighted majority voting converges with infinite samples, and the limit only depends on the distribution modeled by the language model (and the reward model). This theoretical finding is also aligned with our empirical findings shown in Sec. 4.2, which show saturation at high sampling budgets. The proofs are presented in Appendix A.
@@
Theorems 1 & 2 state the convergence of the accuracy with increasing number of samples, indicating that the performance gains of using more samples will saturate for any fixed models. The limit is determined by the likelihood of generating the correct answers through all possible reasoning paths (and the likelihood should be viewed as a weighted sum for weighted majority voting). This motivates us to consider inference algorithms that search for “good” reasoning paths, such as the tree-search-based variants detailed in Sec. 3.1.1 & 3.1.2.
@@ 3.1.1 Monte Carlo Tree Search (MCTS)
Monte Carlo Tree Search (MCTS) has proven effective in domains such as board games where strategic decision-making is required (Silver et al., 2016; 2017; Jones, 2021). Recent work has shown that adapting MCTS to the context of LLMs can enhance the text generation process (Zhang et al., 2023; Zhou et al., 2023; Liu et al., 2024; Choi et al., 2023; Chen et al., 2024a; Tian et al., 2024; Chen et al., 2024a). In this context, MCTS is paired with a value model to score and guide the exploration steps. For additional background, we provide a review of MCTS in Appendix B.
Recent work in MCTS or its variants (e.g., Tree of Thoughts (Yao et al., 2023)) mainly focus on improving the performance (e.g., accuracy) on the studied tasks. However, generic comparisons of MCTS with conventional methods like best-of-$n$ and majority voting in terms of computational budget, measured in generated tokens or processing time are scarce or indicate potentially unfavorable cost-performance tradeoffs. For example, MCTS consumes substantially more resources, often requiring dozens of times more generated tokens than simpler methods. Specifically, a significant portion of the paths in the search tree are used to estimate and select nodes, and these paths do not necessarily become a part of the final candidate solution, although MCTS ensures that the sampled solutions comprise high-quality intermediate steps. In contrast, sampling methods generate multiple solutions in parallel and independently, and all the generated sequences are included in the candidate solutions. However, the intermediate steps in these sequences are not guaranteed to be of high quality, as there is no mechanism for pruning poor steps or exploiting promising ones.
This highlights the need for a new tree search method that can achieve a comparable (or better) performance as MCTS, and that is computationally less costly, with a cost similar to weighted majority voting and best-of-$n$. This motivates our new method, Reward Balanced SEarch (Rebase).